OpenStudy (anonymous):

simplify the expression using the rules for exponents (a^1b^3) ^3/4 / (ab)^-3/4

OpenStudy (anonymous):

For starters, we can see that a^1 is a. Let's start with numerator: (ab^3)^(3/4) A property of exponents is that (ab)^x = a^x * b^x So we get: a^(3/4) * b (9/4) Onto the denominator: Apply the same rule: a^(-3/4) * b^(-3/4) Put the back together: (a^(3/4) * b^(9/4))/ (a^(-3/4) * b^(-3/4) Applying the property that (a^x)/(a^y) = a^(x - y) and the properties of fractions where (ab/cd = a/c * b/d): a^(6/8) * b^(3)